$SPAD/src/input classtalk.input

Timothy Daly December 3, 2016

Abstract

These are examples from the talk “Axiom in an Educational Setting”.

Contents

2

— * —

)spool classtalk.output
)set message test on
)set message auto off
)set break resume
)clear all

———-

1 Numbers

— * —

--S 1 of 72
1
--R
--R   (1)  1
--R                                                        Type: PositiveInteger
--E 1
--S 2 of 72
1/2
--R
--R        1
--R   (2)  -
--R        2
--R                                                      Type: Fraction(Integer)
--E 2
--S 3 of 72
3+4*%i
--R
--R   (3)  3 + 4%i
--R                                                       Type: Complex(Integer)
--E 3
--S 4 of 72
3.4
--R
--R   (4)  3.4
--R                                                                  Type: Float
--E 4
--S 5 of 72
X::ROMAN
--R

--R   (5)  X
--R                                                           Type: RomanNumeral
--E 5
--S 6 of 72
binary(5)
--R
--R   (6)  101
--R                                                        Type: BinaryExpansion
--E 6
--S 7 of 72
factor(60)
--R
--R         2
--R   (7)  2 3 5
--R                                                      Type: Factored(Integer)
--E 7
--S 8 of 72
q:=(y-1)*x*(z+5)
--R
--R   (8)  (x y - x)z + 5x y - 5x
--R                                                    Type: Polynomial(Integer)
--E 8
--S 9 of 72
factor q
--R
--R   (9)  x(y - 1)(z + 5)
--R                                          Type: Factored(Polynomial(Integer))
--E 9
--S 10 of 72
eval(q,[x=5,y=6,z=7])
--R
--R   (10)  300
--R                                                    Type: Polynomial(Integer)
--E 10
--S 11 of 72
eval(q,[x=5,y=6])
--R
--R   (11)  25z + 125
--R                                                    Type: Polynomial(Integer)
--E 11

———-

2 Trigonometry

— * —

--S 12 of 72
b:=[log a, exp a, asin a, acos a, atan a, acot a, sinh a]
--R
--R                   a
--R   (12)  [log(a),%e ,asin(a),acos(a),atan(a),acot(a),sinh(a)]
--R                                              Type: List(Expression(Integer))
--E 12
--S 13 of 72
[exp b.1, log b.2, sin b.3, cos b.4, tan b.5, cot b.6, asinh b.7]
--R
--R   (13)  [a,a,a,a,a,a,a]
--R                                              Type: List(Expression(Integer))
--E 13
--S 14 of 72
a:=.7
--R
--R   (14)  0.7
--R                                                                  Type: Float
--E 14
--S 15 of 72
b:=[log a, exp a, asin a, acos a, atan a, acot a, sinh a]
--R
--R   (15)
--R   [- 0.3566749439 3873237891, 2.0137527074 704765216, 0.7753974966 1075306374,
--R    0.7953988301 8414355549, 0.6107259643 8920861654, 0.9600703624 0568800269,
--R    0.7585837018 3953350346]
--R                                                            Type: List(Float)
--E 15
--S 16 of 72
[exp b.1, log b.2, sin b.3, cos b.4, tan b.5, cot b.6, asinh b.7]
--R
--R   (16)  [0.7,0.7,0.7,0.7,0.7,0.7,0.7]
--R                                                            Type: List(Float)
--E 16
--S 17 of 72
simplify(sin(x)**2+cos(x)**2)
--R
--R   (17)  1
--R                                                    Type: Expression(Integer)
--E 17

———-

3 Polynomial Manipulations

— * —

)clear all
--S 18 of 72
eq1:=A*x^2 + B*x*y + C*y^2 + D*x + E*y + F
--R
--R           2                   2
--R   (1)  C y  + (B x + E)y + A x  + D x + F
--R                                                    Type: Polynomial(Integer)
--E 18
--S 19 of 72
rotatex:=x'*cos(t)-y'*sin(t)
--R
--R   (2)  - y' sin(t) + x' cos(t)
--R                                                    Type: Expression(Integer)
--E 19
--S 20 of 72
rotatey:=x'*sin(t)+y'*cos(t)
--R
--R   (3)  x' sin(t) + y' cos(t)
--R                                                    Type: Expression(Integer)
--E 20
--S 21 of 72
eval(eq1,[x=rotatex, y=rotatey])
--R
--R   (4)
--R          2                 2       2
--R     (A y'  - B x' y' + C x' )sin(t)
--R   +
--R             2                        2
--R     ((- B y'  + (2C - 2A)x' y' + B x' )cos(t) - D y' + E x')sin(t)
--R   +
--R          2                 2       2
--R     (C y'  + B x' y' + A x' )cos(t)  + (E y' + D x')cos(t) + F
--R                                                    Type: Expression(Integer)
--E 21

———-

4 Polynomials over Simple Algebraic Extension Fields

— * —

)clear all
--S 22 of 72
a:=rootOf(a^2+a+1)
--R
--R   (1)  a
--R                                                        Type: AlgebraicNumber
--E 22
--S 23 of 72
factor(x^2+3)
--R
--R         2
--R   (2)  x  + 3
--R                                          Type: Factored(Polynomial(Integer))
--E 23
--S 24 of 72
factor(x^2+3,[a])
--R
--R   (3)  (x - 2a - 1)(x + 2a + 1)
--R                                  Type: Factored(Polynomial(AlgebraicNumber))
--E 24
--S 25 of 72
definingPolynomial(a)
--R
--R         2
--R   (4)  a  + a + 1
--R                                                        Type: AlgebraicNumber
--E 25
--S 26 of 72
zerosOf(b^2+b+1,b)
--R
--R          +---+        +---+
--R         \|- 3  - 1 - \|- 3  - 1
--R   (5)  [----------,------------]
--R              2           2
--R                                              Type: List(Expression(Integer))
--E 26

———-

5 Derivatives

— * —

--S 27 of 72
differentiate(sin(x),x)
--R
--R   (6)  cos(x)
--R                                                    Type: Expression(Integer)
--E 27
--S 28 of 72
differentiate(sin(x),x,2)
--R
--R   (7)  - sin(x)
--R                                                    Type: Expression(Integer)
--E 28
--S 29 of 72
differentiate(cos(z)/(x^2+y^3),[x,y,z],[1,2,3])
--R
--R                    4      3
--R            (- 84x y  + 24x y)sin(z)
--R   (8)  --------------------------------
--R         12     2 9     4 6     6 3    8
--R        y   + 4x y  + 6x y  + 4x y  + x
--R                                                    Type: Expression(Integer)
--E 29
--S 30 of 72
y:=operator y
--R
--R   (9)  y
--R                                                          Type: BasicOperator
--E 30
--S 31 of 72
deqx:=D(y(x),x,2)+D(y(x),x)+y(x)
--R
--R
--R          ,,       ,
--R   (10)  y  (x) + y (x) + y(x)
--R
--R                                                    Type: Expression(Integer)
--E 31
--S 32 of 72
solve(deqx,y,x)
--R

--R                                              x     x
--R                                      +-+   - -   - -      +-+
--R                                    x\|3      2     2    x\|3
--R   (11)  [particular= 0,basis= [cos(-----)%e   ,%e   sin(-----)]]
--R                                      2                    2
--RType: Union(Record(particular: Expression(Integer),basis: List(Expression(Integer))),...)
--E 32

———-

6 Limits

— * —

)clear all
--S 33 of 72
limit((x^2-3*x+2)/(x^2-1),x=1)
--R
--R          1
--R   (1)  - -
--R          2
--R            Type: Union(OrderedCompletion(Fraction(Polynomial(Integer))),...)
--E 33
--S 34 of 72
limit(x*log(x),x=0)
--R
--R   (2)  [leftHandLimit= "failed",rightHandLimit= 0]
--RType: Union(Record(leftHandLimit: Union(OrderedCompletion(Expression(Integer)),"failed"),rightHandLimit: Union(OrderedCompletion(Expression(Integer)),"failed")),...)
--E 34
--S 35 of 72
limit(sinh(a*x)/tan(b*x),x=0)
--R
--R        a
--R   (3)  -
--R        b
--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 35
--S 36 of 72
limit(sqrt(3*x^2+1)/(5*x),x=%plusInfinity)
--R
--R         +-+
--R        \|3
--R   (4)  ----
--R          5

--R                      Type: Union(OrderedCompletion(Expression(Integer)),...)
--E 36
--S 37 of 72
complexLimit((2+z)/(1-z),z=%infinity)
--R
--R   (5)  - 1
--R                      Type: OnePointCompletion(Fraction(Polynomial(Integer)))
--E 37

———-

7 Indefinite Integration

— * —

)clear all
--S 38 of 72
integrate(1+sqrt(x)/x,x)
--R
--R          +-+
--R   (1)  2\|x  + x
--R                                         Type: Union(Expression(Integer),...)
--E 38
--S 39 of 72
integrate(sin(x)/x,x)
--R
--R   (2)  Si(x)
--R                                         Type: Union(Expression(Integer),...)
--E 39

———-

This used to give the answer:

\frac{\sqrt{x}\sqrt{π}erf(x\sqrt{a})}{2a}

— * —

--S 40 of 72
integrate(exp(-a*x^2),x)
--R
--R           x       2
--I         ++    - %Q a
--I   (3)   |   %e      d%Q
--R        ++

--R                                         Type: Union(Expression(Integer),...)
--E 40
--S 41 of 72
integrate(sin(x)/x^2,x)
--R
--R           x
--I         ++  sin(%Q)
--I   (4)   |   ------- d%Q
--R        ++       2
--I               %Q
--R                                         Type: Union(Expression(Integer),...)
--E 41

———-

8 Definite Integration

— * —

)clear all
--S 42 of 72
integrate(exp(-x)/sqrt(x),x=0..%plusInfinity)
--R
--R         _1
--R   (1)  | (-)
--R           2
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 42
--S 43 of 72
integrate(1/x^2,x=-1..1)
--R
--R
--RDaly Bug
--R   >> Error detected within library code:
--R   integrate: pole in path of integration
--R
--R   Continuing to read the file...
--R
--E 43
)clear all

———-

This used to return

\frac{4 log (4) − 8 log (2) + 3π}{12}

— * —

--S 44 of 72
integrate(sin(x)^3/(sin(x)^3+cos(x)^3),x=0..%pi/2,"noPole")
--R
--R        2log(16) - 4log(4) + 3%pi
--R   (1)  -------------------------
--R                    12
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 44
--S 45 of 72
integrate(exp(-x^2)*log(x)^2,x=0..%plusInfinity)
--R
--R         _1             1     _1         1 2
--R        | (-)polygamma(1,-) + | (-)digamma(-)
--R           2             2       2         2
--R   (2)  --------------------------------------
--R                           8
--R                  Type: Union(f1: OrderedCompletion(Expression(Integer)),...)
--E 45

———-

9 Laplace Transformations

— * —

)clear all
--S 46 of 72
laplace(sin(a*t)*cosh(a*t)-cos(a*t)*sinh(a*t),t,s)
--R
--R             3
--R           4a
--R   (1)  --------
--R         4     4
--R        s  + 4a
--R                                                    Type: Expression(Integer)
--E 46
--S 47 of 72
laplace(2/t * (1-cos(a*t)),t,s)
--R
--R             2    2
--R   (2)  log(s  + a ) - 2log(s)
--R                                                    Type: Expression(Integer)

--E 47
--S 48 of 72
laplace((exp(a*t)-exp(b*t))/t,t,s)
--R
--R   (3)  - log(s - a) + log(s - b)
--R                                                    Type: Expression(Integer)
--E 48
--S 49 of 72
laplace(exp(a*t+b)*Ei(c*t),t,s)
--R
--R          b    s + c - a
--R        %e log(---------)
--R                   c
--R   (4)  -----------------
--R              s - a
--R                                                    Type: Expression(Integer)
--E 49

———-

10 Clifford Algebras

CliffordAlgebra(n,K,Q) defines a vector space of dimension 2^{n} over K, given a quadratic form Q on K^{n} (e.q. quaternions).

— * —

)clear all
--S 50 of 72
K:=Fraction Polynomial Integer
--R
--R   (1)  Fraction(Polynomial(Integer))
--R                                                                 Type: Domain
--E 50
--S 51 of 72
qf:QFORM(2,K):=quadraticForm matrix([[-1,0],[0,-1]])$(SQMATRIX(2,K))
--R
--R        +- 1   0 +
--R   (2)  |        |
--R        + 0   - 1+
--R                         Type: QuadraticForm(2,Fraction(Polynomial(Integer)))
--E 51
--S 52 of 72
i:=e(1)$CLIF(2,K,qf)
--R

--R   (3)  e
--R         1
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 52
--S 53 of 72
j:=e(2)$CLIF(2,K,qf)
--R
--R   (4)  e
--R         2
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 53
--S 54 of 72
k:=i*j
--R
--R   (5)  e e
--R         1 2
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 54
--S 55 of 72
x:=a+b*i+c*j+d*k
--R
--R   (6)  a + b e  + c e  + d e e
--R               1      2      1 2
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 55
--S 56 of 72
y:=m+f*i+g*j+h*k
--R
--R   (7)  m + f e  + g e  + h e e
--R               1      2      1 2
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 56
--S 57 of 72
x+y
--R
--R   (8)  m + a + (f + b)e  + (g + c)e  + (h + d)e e
--R                        1           2           1 2
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 57
--S 58 of 72
x*y
--R
--R   (9)
--R     a m - d h - c g - b f + (b m + c h - d g + a f)e

--R                                                     1
--R   +
--R     (c m - b h + a g + d f)e  + (d m + a h + b g - c f)e e
--R                             2                           1 2
--R       Type: CliffordAlgebra(2,Fraction(Polynomial(Integer)),[[-1,0],[0,-1]])
--E 58

———-

11 Taylor Series

— * —

)clear all
--S 59 of 72
taylor(sin(x),x=0)
--R
--R            1  3    1   5     1   7      1    9      11
--R   (1)  x - - x  + --- x  - ---- x  + ------ x  + O(x  )
--R            6      120      5040      362880
--R                        Type: UnivariateTaylorSeries(Expression(Integer),x,0)
--E 59

———-

12 Laurent Series

— * —

--S 60 of 72
laurent(x/log(x),x=1)
--R
--R   (2)
--R            - 1   3    5            1        2    11        3    11         4
--R     (x - 1)    + - + -- (x - 1) - -- (x - 1)  + --- (x - 1)  - ---- (x - 1)
--R                  2   12           24            720            1440
--R   +
--R      271         5    13         6     7297         7     425         8
--R     ----- (x - 1)  - ---- (x - 1)  + ------- (x - 1)  - ------ (x - 1)
--R     60480            4480            3628800            290304
--R   +
--R       530113         9            10
--R     --------- (x - 1)  + O((x - 1)  )
--R     479001600
--R                       Type: UnivariateLaurentSeries(Expression(Integer),x,1)

--E 60

———-

13 Puiseux Series

— * —

--S 61 of 72
puiseux(sqrt(sec(x)),x=3*%pi/2)
--R
--R
--R                    1                3                 7
--R                  - -                -                 -
--R             3%pi   2    1      3%pi 2    1       3%pi 2          3%pi 5
--R   (3)  (x - ----)    + -- (x - ----)  + --- (x - ----)  + O((x - ----) )
--R               2        12        2      160        2               2
--R               Type: UnivariatePuiseuxSeries(Expression(Integer),x,(3*%pi)/2)
--E 61

———-

14 General Series

— * —

--S 62 of 72
series(x^x,x=0)
--R
--R   (4)
--R                         2            3            4            5
--R                   log(x)   2   log(x)   3   log(x)   4   log(x)   5
--R     1 + log(x)x + ------- x  + ------- x  + ------- x  + ------- x
--R                      2            6            24          120
--R   +
--R           6            7            8            9            10
--R     log(x)   6   log(x)   7   log(x)   8   log(x)   9   log(x)    10      11
--R     ------- x  + ------- x  + ------- x  + ------- x  + -------- x   + O(x  )
--R       720          5040        40320        362880       3628800
--R                  Type: GeneralUnivariatePowerSeries(Expression(Integer),x,0)
--E 62

———-

15 Matrices

— * —

)clear all
--S 63 of 72
m:=matrix [[1,2],[3,4]]
--R
--R        +1  2+
--R   (1)  |    |
--R        +3  4+
--R                                                        Type: Matrix(Integer)
--E 63
--S 64 of 72
4*m*(-5)
--R
--R        +- 20  - 40+
--R   (2)  |          |
--R        +- 60  - 80+
--R                                                        Type: Matrix(Integer)
--E 64
--S 65 of 72
n:=matrix [[1,0,-2],[-3,5,1]]
--R
--R        + 1   0  - 2+
--R   (3)  |           |
--R        +- 3  5   1 +
--R                                                        Type: Matrix(Integer)
--E 65
--S 66 of 72
m*n
--R
--R        +- 5  10   0 +
--R   (4)  |            |
--R        +- 9  20  - 2+
--R                                                        Type: Matrix(Integer)
--E 66
--S 67 of 72
hilb:=matrix([[1/(i+j) for i in 1..3] for j in 1..3])
--R
--R        +1  1  1+
--R        |-  -  -|
--R        |2  3  4|
--R        |       |
--R        |1  1  1|

--R   (5)  |-  -  -|
--R        |3  4  5|
--R        |       |
--R        |1  1  1|
--R        |-  -  -|
--R        +4  5  6+
--R                                              Type: Matrix(Fraction(Integer))
--E 67
--S 68 of 72
inverse(hilb)
--R
--R        + 72    - 240   180 +
--R        |                   |
--R   (6)  |- 240   900   - 720|
--R        |                   |
--R        + 180   - 720   600 +
--R                                   Type: Union(Matrix(Fraction(Integer)),...)
--E 68

———-

16 Systems of Equations

— * —

)clear all
--S 69 of 72
solve([x+y+z=8,3*x-2*y+z=0,x+2*y+2*z=17],[x,y,z])
--R
--R   (1)  [[x= - 1,y= 2,z= 7]]
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--E 69
--S 70 of 72
solve([x+2*y+3*z=2,2*x+3*y+4*z=2,3*x+4*y+5*z=2],[x,y,z])
--R
--I   (2)  [[x= %W - 2,y= - 2%W + 2,z= %W]]
--R                    Type: List(List(Equation(Fraction(Polynomial(Integer)))))
--E 70
--S 71 of 72
solve([[1,1,1],[3,-2,1],[1,2,2]],[8,0,17])
--R
--R   (3)  [particular= [- 1,2,7],basis= [[0,0,0]]]
--RType: Record(particular: Union(Vector(Fraction(Integer)),"failed"),basis: List(Vector(Fraction(Integer))))
--E 71

--S 72 of 72
solve([[1,2,3],[2,3,4],[3,4,5]],[2,2,2])
--R
--R   (4)  [particular= [- 2,2,0],basis= [[1,- 2,1]]]
--RType: Record(particular: Union(Vector(Fraction(Integer)),"failed"),basis: List(Vector(Fraction(Integer))))
--E 72
)spool
)lisp (bye)

———-

References^{[1] nothing}

20